sum of 3 prime nos is 100........one of them differ by  36........find the largest no.......EXPLAIN!!!!!

Grade:9

1 Answers

Amitabh Shrivastava
13 years ago

Let the nos be a,b,c

a+b+c=100....[1]

a-b=36....[2]

From [1]and[2]

      2b+c=64

As 2b and 64 are even c must be even.

Now the only even prime is 2. Hence c is 2.

2b+2=64=>b=31

a=36+b=67

the largest no. is 67